40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Coating Ring-down Measurement Lab elog  Not logged in ELOG logo
Message ID: 305     Entry time: Tue Feb 14 10:55:13 2017
Author: Gabriele 
Type: General 
Category: Daily Progress 
Subject: Something fishy with bay 4 of the new setup 

I did two set of measurements with the new coated samples from Montreal. I reshuffled the position in the two measurements. In both cases, the measurement being performed in bay 4 was bad, in the sense that it was very hard to see excited modes. Since the two measurements were carried out with two different disks, it's clear it is a problem with that setup.

SOLVED: there was a connection problem for the DAC output signal controlling the switch

REALLY SOLVED: it was not a cabling issue. The power supply for the switching box had the current limiter on: when all four switches are closed, the box drain about 270mA, which is more than the limit of 250mA. Therefore the power supply voltage dropped and only three switches were actually closed. I switched the power supply to 500mA range and maxed the current limit. Now all four switches are working properly

ELOG V3.1.3-